# Unseen from Seen: Rewriting Observation-Instruction Using Foundation Models for Augmenting Vision-Language Navigation

### v1

@@ Abstract
Data scarcity is a long-standing challenge in the Vision-Language Navigation (VLN) field, which extremely hinders the generalization of agents to unseen environments. Previous works primarily rely on additional simulator data or web-collected images/videos to improve the generalization. However, the simulator environments still face limited diversity, and the web-collected data often requires extensive labor to remove the noise. In this paper, we propose a Rewriting-driven AugMentation (RAM) paradigm for VLN, which directly creates the unseen observation-instruction pairs via rewriting human-annotated training data. Benefiting from our rewriting mechanism, new observation-instruction can be obtained in both simulator-free and labor-saving manners to promote generalization. Specifically, we first introduce Object-Enriched Observation Rewriting, where we combine Vision-Language Models (VLMs) and Large Language Models (LLMs) to derive rewritten object-enriched scene descriptions, enabling observation synthesis with diverse objects and spatial layouts via Text-to-Image Generation Models (T2IMs). Then, we propose Observation-Contrast Instruction Rewriting, which generates observation-aligned rewritten instructions by requiring LLMs to reason the difference between original and new observations. We further develop a mixing-then-focusing training strategy with a random observation cropping scheme, effectively enhancing data distribution diversity while suppressing augmentation data noise during training. Experiments on both the discrete environments (R2R, REVERIE, and R4R datasets) and continuous environments (R2R-CE dataset) show the superior performance and impressive generalization ability of our method. Code is available at https://github.com/SaDil13/VLN-RAM.
Index Terms:
Vision-Language Navigation, Data Augmentation, Foundation Models, Text-to-Image Generation.
@@ I Introduction
Vision-Language Navigation (VLN) [1],[2],[3],[4], as a critical research path to achieve embodied intelligence, defines the task that requires an embodied agent to navigate through complex 3D environments following natural language instructions. To enable VLN agents to learn the observation-instruction alignment knowledge and navigation skills, researchers have provided elaborately annotated trajectory-instruction data from the simulator environments [5] for navigation training. Nevertheless, the limited amount of high-quality manually annotated VLN data, i.e., the data scarcity issue, greatly harms the generalization of existing agents to unseen environments with various possible spatial layouts and cooccurrence objects.
To address the data scarcity problem, prior studies have investigated numerous data augmentation strategies, which can be roughly divided into two categories: simulator-based and web-based. Simulator-based methods [6],[7],[8],[9],[10],[11] rely on the original Matterport3D [5] environment or additional simulators like HM3D [12] and Gibson [13] to perform augmentation trajectory sampling and use the Speaker [7] model for instruction generation. In contrast, web-based approaches[14],[15] collect room images or room-tour videos from the web for trajectory augmentation, and adopt the template-based strategy to generate instructions. Despite the promising performance improvement, these approaches still have the following limitations. For trajectory generation, the simulator data is always constrained in specific environments and hard to extend to various scenarios. On the other hand, web images/videos usually contain extensive noise and demand labor-intensive data cleaning. For instruction augmentation, Speaker-generated instructions are usually uninformative while template-based instructions may have limited flexibility.
@@
Experimental results on R2R [1], REVERIE [2], and R4R [16] show the superior performance of RAM over existing methods especially in unseen scenarios, demonstrating the effectiveness of our proposed rewriting-driven augmentation paradigm in improving generalization. Through only small-scale data augmentation, our RAM achieves comparable or even better performance to a recent state-of-the-art method [10] that relies on large-scale data from additional simulators on R2R and REVERIE. The transfer of our RAM to the R2R-CE [17] dataset further reveals the promising generalization ability of the proposed method.
@@ II-A Vision-Language Navigation
Some recent works have attempted to unleash the rich world knowledge from the foundation models, such as Large Language Models (LLMs) [44],[45] and Vision Language Models (VLMs) [46], [47], to improve the generalization ability of the VLN agent [48],[49],[50],[51]. For example, NavGPT [48] builds a purely GPT4-based navigation agent that reasons the action decision based on the textual represented visual observations and navigational histories. DiscussNav [49] incorporates multiple foundation models like InstructBLIP [47] and GPT4 [45] to address different navigation inputs and produce comprehensive reasoning for decision.
@@ II-C LLM-Driven Robotic Data Generation
In this work, we harness the physical world knowledge, reasoning ability, and language skills of the LLM, and integrate the LLM with other foundation models like T2IMs for conducting observation-instruction rewriting for VLN. As a result, our rewritten observations will have reasonable and diverse objects conforming to real-world scenarios. Moreover, the rewritten instructions will be more flexible and informative compared to Speaker-generated instructions and template-based instructions in previous VLN approaches.
@@ III-B Object-Enriched Observation Rewriting
Object-Enriched Observation Rewriting consists of two primary procedures, namely Object-Enriched Scene Description Rewriting and Panorama-to-View Observation Generation. During Object-Enriched Scene Description Rewriting, we utilize VLMs to collect scene description $C_{t}$ for original observation $O_{t}$, then we leverage LLMs to provide object cooccurrence knowledge and generate rewritten object-enriched scene descriptions $C_{t}^{r}$. In Panorama-to-View Observation Generation, we obtain rewritten observations $O_{t}^{r}$ by feeding $C_{t}^{r}$ to T2IMs to directly synthesize rewritten panoramas, followed by a simple panorama discretization algorithm to produce new single-view observations.
@@
To constrain the LLM to explicitly indicate possible scene objects in its generated scene descriptions, we ask the LLM in $P^{c}$ to generate both newly added objects $\{B_{t,n}\}_{n=1}^{N}$ and rewritten scene descriptions $C^{r}_{t}$ by feeding an in-context task example, as shown in Fig. 3. Additionally, to encourage the subsequent generated observations via T2IMs based on the rewritten scene descriptions $C^{r}_{t}$ to indicate different salient objects, which can enhance the diversity potentially, we also require the LLM in $P^{c}$ to change the representation of original scene descriptions for highlighting different objects. The reason why we ask the LLM to explicitly specify objects added to the synthetic scene is to facilitate adding objects with specific numbers as well as manually checking the generation quality of rewritten scene descriptions. And we do not utilize the outputs of added objects in the subsequent stages of our method. Therefore, with our scene description rewriting prompt $P^{c}$ and the original scene description $C_{t}$, we obtain the rewritten object-enriched scene descriptions $C^{r}_{t}$ accompanied by added objects $\{B_{t,n}\}_{n=1}^{N}$ by:
@@
For panorama generation, we choose MultiDiffusion [60], which is capable of generating images with desired aspect ratios and informative contents to generate the panoramic images. We employ the Equirec2Perspec algorithm111 https://github.com/fuenwang/Equirec2Perspec to discretize a panoramic image into 36 perspective images following the same setting in VLN [1]. Initially, the Equirec2Perspec algorithm utilizes the camera parameters, encompassing the Field of View (FOV), heading ($\theta$), and elevation ($\phi$), to calculate the inverse matrix of the camera’s projection matrix, denoted as $J_{inv}$, and the rotation matrix, denoted as $R$:
$\begin{split}J_{inv}&=\mathrm{MC_{proj}}(\mathrm{FOV},\theta,\phi),\\ R&=\mathrm{MC_{rot}}(\theta,\phi),\\ \end{split}$ (3)
where $\mathrm{MC_{proj}}(\cdot)$ is the camera’s projection matrix calculation function and $\mathrm{MC_{rot}}(\cdot)$ is the roration matrix calculation function.
@@
[…]
@@ III-C Observation-Contrast Instruction Rewriting
Specifically, we first perform Sequential Landmark Grounding, i.e., obtaining the grounded landmark $U_{t}$ for each observation $O_{t}$ in the original trajectory from the original instruction $I$. Then, we conduct New Observation Description Collection to derive descriptions $C^{\prime}_{t}$ for the rewritten observations $O^{r}_{t}$. Finally, we obtain rewritten instructions $I^{r}$ by asking the LLM to: 1) contrast each $U_{t}$ and $C^{\prime}_{t}$ to replace $U_{t}$ in $I$ with the landmark appearing in $C^{\prime}_{t}$ at the appropriate timestep, and 2) change the original actional representations in $I$ to their synonyms. Consequently, our rewritten instructions can be observation-aligned with flexible representations. Moreover, benefiting from rewriting human-annotated instructions, the generated instructions can be essentially more informative.
@@
${U}_{t}=\mathop{\mathrm{argmax}}\limits_{U_{k}}\mathrm{Sim}\left(F^{v}\left(G_{t}\right),F^{t}\left(U_{k}\right)\right),\vspace{-0.2cm}$ (5)
@@
Instruction Rewriting by Observation Contrast. As shown in Fig. 3, we build the instruction rewriting prompt $P^{i}$ with the task description of “Rewrite the original instruction by replacing the object (scene) in the original instruction with the one in the new observation…”, followed by an in-context task example, which teaches the LLM to perform observation contrast as well as change the representations of actional descriptions to produce rewritten instructions. Therefore, with our instruction rewriting prompt $P^{i}$, the grounded landmark information $\{U_{t}\}_{t=1}^{T}$, the new observation descriptions $\{C^{\prime}_{t}\}_{t=1}^{T}$, and the original instruction $I$, the rewritten instruction $I^{r}$ is acquired by:
@@ IV-A1 Datasets
[…]
@@ IV-A2 Evaluation Metrics
TABLE II: Comparison with existing methods on R2R. *: 900 extra simulator scenes for training and using CLIP ViT H/14 as image features. †: using CLIP ViT B/16 as image features. ‡: using CLIP ViT L/14 as image features.
@@ IV-A3 Implementation Details
We plug our RAM into a powerful pretraining-based baseline [39]. In the pretraining phase, the agent is trained with three proxy tasks: Masked Language Modeling (MLM), Instruction and Trajectory Matching (ITM), and Single Action Prediction (SAP). In the finetuning phase, the agent is trained with DAGGER [67] method. We add our rewritten data during both the pretraining and finetuning phases to boost the navigation training. The pretraining is conducted for 20k iterations with a batch size of 128 and learning rate of 5e-5 on four NVIDIA RTX 3090 GPUs. Our mixing-then-focusing training strategy is implemented during the finetuning phase, where the maximum training iterations, the batch size, and the learning rate for two stages are set as 20k, 8, and 1e-5, respectively. We resume the best model in stage 1 to train for stage 2. The finetuning is implemented on a single NVIDIA RTX 3090 GPU. The pre-training phase lasts about 50 hours and the fine-tuning phase lasts about 20 hours.
We adopt two kinds of visual encoders, i.e., CLIP ViT B/16 and CLIP ViT L/14 [46] to validate the effectiveness of our method. Like previous works [38],[39],[53], we also add the prevalent data [37] during training. We use Tag2Text [68] as the VLM for observation description collection, and employ ChatGPT [44](with version gpt-3.5-turbo-1106) as the LLM for scene description rewriting and instruction rewriting. For an original trajectory-instruction pair, we collect 3 different trajectories containing new panoramas, and for each new trajectory we obtain one corresponding augmentation instruction. The LLMs, VLMs, T2IMs used in the experiments are all in a zero-shot manner without the need for training consumption. The parameter size of our model is 688.51M.
@@ IV-B Comparison with Existing Approaches
R2R. Table II shows the performance comparison of different methods on R2R, where we can find that our RAM using CLIP ViT L/14 features outperforms all previous works that do not introduce large-scale additional data for training like ScaleVLN [10]. Although ScaleVLN outperforms previous works as well as RAM, it resorts to significantly larger realistic data resources ($\times$352 larger than the R2R dataset) while our RAM only introduces a small scale of generated data ($\times$3 larger than the R2R dataset). Moreover, our RAM with both two kinds of CLIP features outperforms ScaleVLN [10] in Val Seen split. We can also observe that our RAM outperforms baseline DUET [39] especially in Val Unseen when using both CLIP ViT B/16 and L/14 as image features, revealing its superior generalization ability.
REVERIE. Table III compares different methods regarding the navigation and object grounding performance on REVERIE, where we can find that RAM surpasses the baseline [39] largely on both navigation and object grounding metrics under the same visual features on Val Unseen, e.g., the improvements for SR and RGS are $\sim$3.1and $\sim$2.2%, respectively. RAM also significantly outperforms previous works including AutoVLN [9] and ScaleVLN [10] that introduce large-scale additional simulator data on Test Unseen regarding both navigation and object grounding metrics. These results show that our RAM effectively improves the generalization without the need to introduce large-scale additional realistic data.
TABLE IV: Performance comparison on R4R dataset. ‡: using CLIP ViT L/14 as image features.
@@
[…]

### v2

@@ Abstract
Data scarcity is a long-standing challenge in the Vision-Language Navigation (VLN) field, which extremely hinders the generalization of agents to unseen environments. Previous works primarily rely on additional simulator data or web-collected images/videos to improve the generalization. However, the simulator environments still face limited diversity, and the web-collected data often requires extensive labor to remove the noise. In this paper, we propose a Rewriting-driven AugMentation (RAM) paradigm for VLN, which directly creates the unseen observation-instruction pairs via rewriting human-annotated training data. Benefiting from our rewriting mechanism, new observation-instruction pairs can be obtained in both simulator-free and labor-saving manners to promote generalization. Specifically, we first introduce Object-Enriched Observation Rewriting, where we combine Vision-Language Models (VLMs) and Large Language Models (LLMs) to derive rewritten object-enriched scene descriptions, enabling observation synthesis with diverse objects and spatial layouts via Text-to-Image Generation Models (T2IMs). Then, we propose Observation-Contrast Instruction Rewriting, which generates observation-aligned rewritten instructions by requiring LLMs to reason the difference between original and new observations. We further develop a mixing-then-focusing training strategy with a random observation cropping scheme, effectively enhancing data distribution diversity while suppressing augmentation data noise during training. Experiments on both the discrete environments (R2R, REVERIE, and R4R datasets) and continuous environments (R2R-CE dataset) show the superior performance and impressive generalization ability of our method.
@@ I Introduction
Vision-Language Navigation (VLN) [1],[2],[3],[4], as a critical research path to achieve embodied intelligence, defines the task that requires an embodied agent to navigate through complex 3D environments following natural language instructions. To enable VLN agents to learn the observation-instruction alignment knowledge and navigation skills, researchers have provided elaborately annotated trajectory-instruction data from the simulator environments [5] for navigation training. Nevertheless, the limited amount of high-quality manually annotated VLN data, i.e., the data scarcity issue, greatly harms the generalization of existing agents to unseen environments with various possible spatial layouts and co-occurrence objects.
To address the data scarcity problem, prior studies have investigated numerous data augmentation strategies, which can be roughly divided into two categories: simulator-based and web-based. Simulator-based methods [6],[7],[8],[9],[10],[11] rely on the original Matterport3D [5] environment or additional simulators like HM3D [12] and Gibson [13] to perform augmentation trajectory sampling and use the Speaker [7] model for instruction generation. In contrast, web-based approaches[14], [15], collect room images or room-tour videos from the web for trajectory augmentation, and adopt the template-based strategy to generate instructions. Despite the promising performance improvement, these approaches still have the following limitations. For trajectory generation, the simulator data is always constrained in specific environments and hard to extend to various scenarios. On the other hand, web images/videos usually contain extensive noise and demand labor-intensive data cleaning. For instruction augmentation, Speaker-generated instructions are usually uninformative while template-based instructions may have limited flexibility.
@@
Experimental results on R2R [1], REVERIE [2], and R4R [16] show the superior performance of RAM over existing methods especially in unseen scenarios, demonstrating the effectiveness of our proposed rewriting-driven augmentation paradigm in improving generalization. Through only small-scale data augmentation, our RAM achieves comparable or even better performance to a recent State-Of-The-Art (SOTA) method [10] that relies on large-scale data from additional simulators on R2R and REVERIE. The transfer of our RAM to the R2R-CE [17] dataset further reveals the promising generalization ability of the proposed method.
@@ II-A Vision-Language Navigation
Some recent works have attempted to unleash the rich world knowledge from the foundation models, such as Large Language Models (LLMs) [49],[50] and Vision Language Models (VLMs) [51], [52], to improve the generalization ability of the VLN agent [53],[54],[55],[56]. For example, NavGPT [53] builds a purely GPT-4-based navigation agent that reasons the action decision based on the textual represented visual observations and navigational histories. DiscussNav [54] incorporates multiple foundation models like InstructBLIP [52] and GPT-4 [50] to address different navigation inputs and produce comprehensive reasoning for decision.
@@ II-C LLM-Driven Robotic Data Generation
In this work, we harness the physical world knowledge, reasoning ability, and language skills of the LLMs, and integrate the LLMs with other foundation models like T2IMs for conducting observation-instruction rewriting for VLN. As a result, our rewritten observations will have reasonable and diverse objects conforming to real-world scenarios. Moreover, the rewritten instructions will be more flexible and informative compared to Speaker-generated instructions and template-based instructions in previous VLN approaches.
@@ III-B Object-Enriched Observation Rewriting
Object-Enriched Observation Rewriting consists of two primary procedures, namely Object-Enriched Scene Description Rewriting and Panorama-to-View Observation Generation. During Object-Enriched Scene Description Rewriting, we utilize VLMs to collect scene description $C_{t}$ for original observation $O_{t}$, then we leverage LLMs to provide object co-occurrence knowledge and generate rewritten object-enriched scene descriptions $C_{t}^{r}$. In Panorama-to-View Observation Generation, we obtain rewritten observations $O_{t}^{r}$ by feeding $C_{t}^{r}$ to T2IMs to directly synthesize rewritten panoramas, followed by a simple panorama discretization algorithm to produce new single-view observations.
@@
To constrain the LLMs to explicitly indicate possible scene objects in its generated scene descriptions, we ask the LLMs in $P^{c}$ to generate both newly added objects $\{B_{t,n}\}_{n=1}^{N}$ and rewritten scene descriptions $C^{r}_{t}$ by feeding an in-context task example, as shown in Fig. 3. Additionally, to encourage the subsequent generated observations via T2IMs based on the rewritten scene descriptions $C^{r}_{t}$ to indicate different salient objects, which can enhance the diversity potentially, we also require the LLMs in $P^{c}$ to change the representation of original scene descriptions for highlighting different objects. The reason why we ask the LLMs to explicitly specify objects added to the synthetic scene is to facilitate adding objects with specific numbers as well as manually checking the generation quality of rewritten scene descriptions. And we do not utilize the outputs of added objects in the subsequent stages of our method. Therefore, with our scene description rewriting prompt $P^{c}$ and the original scene description $C_{t}$, we obtain the rewritten object-enriched scene descriptions $C^{r}_{t}$ accompanied by added objects $\{B_{t,n}\}_{n=1}^{N}$ by:
@@
For panorama generation, we choose MultiDiffusion [65], which is capable of generating images with desired aspect ratios and informative contents to generate the panoramic images. We employ the Equirec2Perspec algorithm [^1] https://github.com/fuenwang/Equirec2Perspec to discretize a panoramic image into 36 perspective images following the same setting in VLN [1]. Initially, the Equirec2Perspec algorithm utilizes the camera parameters, encompassing the Field of View (FOV), heading ($\theta$), and elevation ($\phi$), to calculate the inverse matrix of the camera’s projection matrix, denoted as $J_{inv}$, and the rotation matrix, denoted as $R$:
[…]
where $\mathrm{MC_{proj}}(\cdot)$ is the camera’s projection matrix calculation function and $\mathrm{MC_{rot}}(\cdot)$ is the rotation matrix calculation function.
@@
where the $\mathrm{Equirec2Perspec}$ function is the synthesis of a series of coordinate conversion functions. As shown in Fig. 2, the generated single-view images (see “New GT Actions”) have a similar field of view effect to the original single-view images in the Matterport3D [5] simulator.
@@ III-C Observation-Contrast Instruction Rewriting
Specifically, we first perform Sequential Landmark Grounding, i.e., obtaining the grounded landmark $U_{t}$ for each observation $O_{t}$ in the original trajectory from the original instruction $I$. Then, we conduct New Observation Description Collection to derive descriptions $C^{\prime}_{t}$ for the rewritten observations $O^{r}_{t}$. Finally, we obtain rewritten instructions $I^{r}$ by asking the LLMs to: 1) contrast each $U_{t}$ and $C^{\prime}_{t}$ to replace $U_{t}$ in $I$ with the landmark appearing in $C^{\prime}_{t}$ at the appropriate timestep, and 2) change the original actional representations in $I$ to their synonyms. Consequently, our rewritten instructions can be observation-aligned with flexible representations. Moreover, benefiting from rewriting human-annotated instructions, the generated instructions can be essentially more informative.
@@
${U}_{t}=\mathop{\mathrm{argmax}}\limits_{U_{k}}\mathrm{Sim}\left(F^{v}\left(G_{t}\right),F^{t}\left(U_{k}\right)\right),\vskip-5.69046pt$ (5)
@@
Instruction Rewriting by Observation Contrast. As shown in Fig. 3, we build the instruction rewriting prompt $P^{i}$ with the task description of “Rewrite the original instruction by replacing the object (scene) in the original instruction with the one in the new observation…”, followed by an in-context task example, which teaches the LLMs to perform observation contrast as well as change the representations of actional descriptions to produce rewritten instructions. Therefore, with our instruction rewriting prompt $P^{i}$, the grounded landmark information $\{U_{t}\}_{t=1}^{T}$, the new observation descriptions $\{C^{\prime}_{t}\}_{t=1}^{T}$, and the original instruction $I$, the rewritten instruction $I^{r}$ is acquired by:
@@ IV-A1 Datasets
We mainly evaluate our RAM on three discrete VLN benchmarks: R2R [1], REVERIE [2], and R4R [16]. R2R contains 90 indoor scenes with 7189 trajectories. Each trajectory is accompanied by 3 step-by-step instructions. On average, an instruction contains 32 words, and each ground-truth path is formed by seven nodes with a total length of 10mm. REVERIE replaces the fine-grained instructions in R2R with high-level instructions that are only related to the target position and object. R4R creates longer instructions and trajectories by concatenating two adjacent tail-to-head trajectories in R2R. Both R2R and REVERIE split the evaluation set into Val Seen, Val Unseen, and Test Unseen, while R4R only splits the evaluation set into Val Seen and Val Unseen. “Val Seen” means new trajectory-instruction pair data but collected from the same houses as those in the training dataset. On the other hand, “Val Unseen” and “Test Unseen” stand for entirely different VLN data collected from additional houses which are different from the houses in the training dataset.
@@ IV-A2 Evaluation Metrics
TABLE II: Comparison with existing methods on R2R. *: 900 extra simulator scenes for training and using CLIP ViT-H/14 as image features. †: using CLIP ViT-B/16 as image features. ‡: using CLIP ViT-L/14 as image features.
@@ IV-A3 Implementation Details
We plug our RAM into a powerful pretraining-based baseline [39]. In the pretraining phase, the agent is trained with three proxy tasks: Masked Language Modeling (MLM), Instruction and Trajectory Matching (ITM), and Single Action Prediction (SAP). In the finetuning phase, the agent is trained with DAGGER [71] method. We add our rewritten data during both the pretraining and finetuning phases to boost the navigation training. The pretraining is conducted for 20k iterations with a batch size of 128 and learning rate of 5e-5 on four NVIDIA RTX 3090 GPUs. Our mixing-then-focusing training strategy is implemented during the finetuning phase, where the maximum training iterations, the batch size, and the learning rate for two stages are set as 20k, 8, and 1e-5, respectively. We resume the best model in stage 1 to train for stage 2. The finetuning is implemented on a single NVIDIA RTX 3090 GPU. The pretraining phase lasts about 50 hours and the finetuning phase lasts about 20 hours.
We adopt two kinds of visual encoders, i.e., CLIP ViT-B/16 and CLIP ViT-L/14 [51] to validate the effectiveness of our method. Like previous works [38],[39],[58], we also add the prevalent data [36] during training. We use Tag2Text [72] as the VLM for observation description collection, and employ ChatGPT [49](with version gpt-3.5-turbo-1106) as the LLM for scene description rewriting and instruction rewriting. For an original trajectory-instruction pair, we collect 3 different trajectories containing new panoramas, and for each new trajectory we obtain one corresponding augmentation instruction. The LLMs, VLMs, T2IMs used in the experiments are all in a zero-shot manner without the need for training consumption. The parameter size of our model is 688.51M.
@@ IV-B Comparison with Existing Approaches
R2R. Table II shows the performance comparison of different methods on R2R, where we can find that our RAM using CLIP ViT-L/14 features outperforms all previous works that do not introduce large-scale additional data for training like ScaleVLN [10]. Although ScaleVLN outperforms previous works as well as RAM, it resorts to significantly larger realistic data resources ($\times$352 larger than the R2R dataset) while our RAM only introduces a small scale of generated data ($\times$3 larger than the R2R dataset). Moreover, our RAM with both two kinds of CLIP features outperforms ScaleVLN [10] in Val Seen split. We can also observe that our RAM outperforms baseline DUET [39] especially in Val Unseen when using both CLIP ViT-B/16 and ViT-L/14 as image features, revealing its superior generalization ability.
REVERIE. Table III compares different methods regarding the navigation and object grounding performance on REVERIE, where we can find that RAM surpasses the baseline [39] largely on both navigation and object grounding metrics under the same visual features on Val Unseen, e.g., the improvements for SR and RGS are $\sim$3.1% and $\sim$2.2%, respectively. RAM also significantly outperforms previous works including AutoVLN [9] and ScaleVLN [10] that introduce large-scale additional simulator data on Test Unseen regarding both navigation and object grounding metrics. These results show that our RAM effectively improves the generalization without the need to introduce large-scale additional realistic data.
TABLE IV: Performance comparison on R4R dataset. ‡: using CLIP ViT-L/14 as image features.
@@
TABLE V: Results on R2R-CE dataset. $\dagger$}: using the same waypoint predictor [64] for a fair comparison.